# AI-powered simulation-based inference of a genuinely spatial-stochastic model of early mouse embryogenesis Michael A. Ramirez-Sierra1,2, Thomas R. Sokolowski1*

### 23 Feb 2024 v1

@@ Abstract
Understanding how multicellular organisms reliably orchestrate cell-fate decisions is a central challenge in developmental biology. This is particularly intriguing in early mammalian development, where early cell-lineage differentiation arises from processes that initially appear cell-autonomous but later materialize reliably at the tissue level. In this study, we develop a multi-scale, spatial-stochastic simulator of mouse embryogenesis, focusing on inner-cell mass (ICM) differentiation in the blastocyst stage. Our model features biophysically realistic regulatory interactions and accounts for the innate stochasticity of the biological processes driving cell-fate decisions at the cellular scale. We advance event-driven simulation techniques to incorporate relevant tissue-scale phenomena and integrate them with Simulation-Based Inference (SBI), building on a recent AI-based parameter learning method: the Sequential Neural Posterior Estimation (SNPE) algorithm. Using this framework, we carry out a large-scale Bayesian inferential analysis and determine parameter sets that reproduce the experimentally observed system behavior. We elucidate how autocrine and paracrine feedbacks via the signaling protein FGF4 orchestrate the inherently stochastic expression of fate-specifying genes at the cellular level into reproducible ICM patterning at the tissue scale. This mechanism is remarkably independent of the system size. FGF4 not only ensures correct cell lineage ratios in the ICM, but also enhances its resilience to perturbations. Intriguingly, we find that high variability in intracellular initial conditions does not compromise, but rather enhance the accuracy and precision of tissue-level dynamics. Our work provides a genuinely spatial-stochastic description of the biochemical processes driving ICM differentiation and the necessary conditions under which it can proceed robustly.
@@ Introduction
In mammalian development, the cell signaling and fate specification dynamics during the preimplantation stage of mouse embryogenesis serve as key catalysts for understanding this core paradigm. These processes have been extensively investigated through both experimental and theoretical approaches [20, 74, 109, 77, 141, 142, 62, 8, 65, 87, 93, 14, 132, 108, 83, 78], and rely on dynamic cross-interactions between gene expression, molecular signaling at the tissue level, and mechanical cues essential for tissue remodeling [153, 104, 152, 112, 110, 113, 122, 4, 84, 23, 119, 94, 46, 21, 150].
@@
The preimplantation stage in mouse embryos involves two pivotal events transforming the zygote into a blastocyst composed of three distinct cell types [109, 15, 4, 122, 117]. Initially, cell segregation results in the formation of the trophectoderm (TE), which contributes to the embryonic part of the placenta, and the inner cell mass (ICM), an uncommitted bipotent progenitor tissue. Subsequently, a second cell-fate decision within the ICM leads to the differentiation of epiblast (EPI) and primitive endoderm (PRE) tissues. The EPI, marked primarily by NANOG expression, is pluripotent and gives rise to all embryo-proper tissues. In contrast, the PRE, marked by GATA6 expression, contributes to the formation of extraembryonic supporting tissues. Notably, EPI cells secrete FGF4, a signaling ligand that promotes PRE cell specification. The FGF/ERK signaling pathway therefore plays a crucial role in regulating the balance between EPI and PRE cells. Interestingly, FGF4 acts within an external feedback loop that ultimately inhibits its own production, downregulating it in PRE cells [118, 14, 113]. The mechanisms underlying the precise and timely distribution of FGF4 across the ICM, critical for appropriate cell lineage specification, remain elusive [106, 107].
@@
For successful embryogenesis, the TE, EPI, and PRE lineages in the blastocyst must be segregated in specific proportions and within a narrow developmental time window [14, 4]. This specification occurs over approximately 48 hours between embryonic day (E) 2.5 and E4.5 [113, 14]. By E3.0, when blastocyst formation begins, ICM cells co-express NANOG and GATA6. In the further course, EPI and PRE cells emerge asynchronously from ICM cells in a spatially stochastic manner, influenced by the short-range FGF4 signal and regulated by auto- and paracrine feedback loops [73, 4, 106]. A mutually exclusive gene expression profile emerges, with EPI cells exhibiting high NANOG and low GATA6 levels, and vice versa in PRE cells. This process is driven by a gene-regulatory network characterized by self-activation and mutual repression of NANOG and GATA6, and refined by the external feedback via FGF4. The spatio-temporal segregation of EPI and PRE lineages is subsequently coordinated through a mechanical cell-sorting process [150].
@@
Experimental studies have sought to characterize these dynamics and its biophysical parameters, focusing on internal and external developmental cues, molecular mechanisms underpinning gene transcription and mRNA translation, and key signaling pathway components [96, 95, 62, 137, 73]. Nevertheless, this remains challenging, as the involved biochemical species are present at low mRNA and protein abundance, leading to significant biological noise. Here biological noise is categorized as either intrinsic (related to the discrete and stochastic nature of biochemical reactions and molecular transport within cells) or extrinsic (referring to external fluctuations affecting all cells). This noise strongly impacts the reliability of developmental dynamics, particularly during early development [56, 139, 79, 80, 145], and hampers experimental measurements at the same time.
@@
In this study, we investigate cell specification during blastocyst formation in the early mouse embryo under truly stochastic conditions. In order to elucidate mechanisms that enable robust and replicable cell-type proportioning in the ICM, we have constructed a spatial model that allows for exact simulation of the stochastic dynamics of key biochemical species in the developing ICM. This model incorporates three distinct diffusive-signaling modes between neighboring cells: autocrine, paracrine, and membrane-level ligand-exchange (akin to juxtacrine) signaling via FGF4.
@@ Results
In computational modeling of complex biological systems, inferring parameter sets that reproduce experimental observations is a key challenge. This usually requires specific domain knowledge, as the selection of an appropriate parameter search method is chiefly influenced by the particular properties of the problem at hand. One significant hurdle in developmental biology is the lack of a versatile, general-purpose inference method suitable for high-dimensional stochastic models, which typically require large and comprehensive datasets with fine-grained resolution. This expands both in terms of the features measured (e.g., gene expression levels at the single-cell scale rather than bulk measurements) and the temporal resolution (e.g., time-series data capturing dynamics at all relevant scales). However, most experimental studies are unable to simultaneously measure all system variables required for such in-depth mechanistic representation. To bridge this gap, Simulation-Based Inference (SBI) has emerged as a powerful strategy, addressing several of these inferential challenges [27]. Particularly advantageously, SBI allows to learn multidimensional parameter sets that comply with a prescribed behavior formulated in terms of lower-dimensional utility functions, thus mitigating the limitations posed by the scarcity of detailed quantitative data.
@@
Estimate the maximum-a-posteriori (MAP) parameter value that aligns with the desired system behavior represented by the target score time series, using the posteriors predicted by the ANN for the given target score.
@@
For a more detailed description of these steps, see Fig 1 and Model parameter inference framework. In a companion study (in preparation), we compare this SNPE-guided methodology against a strategy inspired by a classical optimization algorithm, specifically simulated annealing; we find that the AI-based approach surpasses its classical counterpart in performance, assuming equal allocation of computational resources. This superiority can be traced back to the innate flexibility of the ANN and the absence of a native parameter interpolation procedure in simulated annealing.
@@
Secondly, blastocyst formation occurs within approximately 1.5-2 days of development [104, 112, 4], as the EPI and PRE populations must reach the required fate proportions 8 to 12 hours before the end of the preimplantation period. This requirement serves as another optimization criterion in our model, postulating that regardless of the underlying differentiation mechanism, the system must attain the target fate proportions by the 40-hour mark.
@@
Fig 2 summarizes the key differences between the ITWT and RTM systems, contrasting the respective inferred parameter values, i.e. the maximum-a-posteriori-probability (MAP) estimates, for both the central intracellular GRN components (top row) and the extracellular signaling topology (bottom row). The most notable difference between the ITWT and RTM models is the reversal of the relationship between the self-activation thresholds of Nanog and Gata6. For the ITWT, the half-saturation threshold of Nanog self-activation (Nanog_NANOG) is lower/stronger than the half-saturation threshold of Gata6 self-activation (Gata6_GATA6). However, both models exhibit a well-balanced mutual repression between Nanog and Gata6, albeit with varying interaction strengths. We present a complete overview of the full model parameter distributions inferred in this study, and a first approximation of the model parameter sensitivities, in the Supporting information, Fig 14 and 15.
@@
The TM-A model displays a decreased accuracy of the ICM specification mechanism due to the absence of self-regulation, though the overall precision of the system remains unaffected (Fig 6[B, D]). Adjustments in other signaling components could potentially correct for this, but also would reduce the system’s ability to buffer against dynamic signaling perturbations.
@@
A central challenge in developmental biology is the precise characterization of spatial patterns, such as the “salt-and-pepper” arrangement reported for the ICM, which has often been described informally in the literature [15, 34, 142, 20]. The term typically implies a random distribution, yet randomness in a mathematical context can take various forms. Recent studies have endeavored to rigorously define this pattern using experimental and theoretical approaches [78, 46, 48, 47]. Here we understand the “salt-and-pepper” pattern as an archetype in which each individual cell-fate decision is independent of the cell fates of its neighbors, which implies a multinomial distribution of cell fates in every tissue neighborhood.
@@
Our approach was guided by two aspects: firstly, the broad inferred posterior distributions of the parameters governing the ICDs, and their moderate sensitivity to value changes (Fig 15[D, H]); secondly, similar assessments that were carried out in previous models of mouse-blastocyst, where mainly the ICD variance was modulated [141, 132]. Taking this into account, we generated 1000 stochastic trajectories for 10 distinct ICDs, modifying their variance from 0% to 200% compared with the baseline value (see Fig 8[A] for examples); the details of ICD modulation are described in Computational experiments. We analyzed both the mean cell count (accuracy) and the corresponding standard deviation (precision) for each cell fate (Fig 8[C, D]).
@@
We quantified the deviations from the typical ICD behavior using the relative fate error (FE), which measures the discrepancy in ICM specification accuracy for each cell fate (UND/EPI/PRE) between the perturbed scenarios and the reference (0% ICP) scenario at 48 hours.
@@
Temporal modulation of FGF4 concentration and the corresponding shift in cell plasticity are key aspects of ICM cell differentiation, with numerous studies documenting their influence [14, 105, 106, 151, 4]. To examine whether our model can replicate the experimentally observed time-dependent responsiveness to FGF4 level changes, we introduced controlled perturbations of FGF4 in our simulations. This involved adding extra FGF4 molecules to each simulated cell at specific time points, simulating the effect of exogenous FGF4.
We assessed the response to FGF4 perturbations in two models (ITWT and TM-FGF4) and at two distinct final simulation times, 48 and 96 hours; the second time point, while biologically irrelevant, was introduced to test whether manipulation of the FGF levels can alter the typical time scale on which cell-fate commitment converges. Exogenous FGF4 addition was carried out at predetermined intervals: hours, with the amount of added FGF4 molecules determined by a mix of Poissonian and binomial distributions (details in Computational experiments). Note that the TM-FGF4 system represents a theoretical mutant with blocked FGF4 production, mimicking a full loss-of-function phenotype for the Fgf4 gene; this means that any FGF4 stems in these systems from the external addition.
Our analysis at the 48-hour mark shows that the TM-FGF4 model is highly responsive to exogenous FGF4 added at the start of the simulation (addition time = 0 hours). Depending on the FGF4 count, it is possible to manipulate the lineage ratios and rescue the PRE fate (Fig 10[E]), which is in line with recent in vitro experiments [106]. However, we identify an end-of-plasticity time point around 24 hours, after which the system becomes insensitive to additional FGF4 and locks into its pre-existing cell lineage proportions (Fig 10[F-H]). In the ITWT system, cell fate ratios can also be manipulated by varying the FGF4 amount, although no distinct end-of-plasticity time point is observed in this model (Fig 10[A-D]).
@@ Discussion
Several phenomenological models have been proposed for blastocyst formation in the mouse embryo [15, 141, 113, 132], but they are primarily deterministic in nature, and as such do not allow for a rigorous investigation of the implications of noise emerging from the basic processes driving this developmental process, neither for quantification of its robustness. To correctly capture the inherent randomness in ICM differentiation, we developed a biophysics-rooted spatial-stochastic model simulated via the Reaction-Diffusion Master Equation (RDME) formalism, and embedded it into a Simulation-Based Inference (SBI) framework building on recent advancements in Machine Learning (ML). Our multi-cellular model mechanistically describes ICM patterning dynamics and its accuracy using biologically realistic lifetimes for the involved biochemical species, providing a biophysically realistic implementation of the mesoscopic processes generating noise at the cell level. Using this combined framework, we inferred multiple parameter distributions that inform our model both in wild-type-like and several mutant-like conditions.
@@
Our inferred theoretical wild-type (ITWT) model recovers key experimental findings, specifically: high reproducibility of the EPI and PRE lineage proportions, the observed timescale of blastocyst formation, amounting to 1.5-2 days of embryonic development, and the indispensability of FGF4-mediated signaling for appropriate ICM patterning [15, 114, 4]. Without FGF4 coupling, the blastocyst fails to establish correct cell fate proportions, displaying a strong bias towards the EPI fate. We thus argue that, given a default naive pluripotent state in ICM-like systems, this cell-fate decision process necessitates a tissue-level coordinating mechanism. Previous experimental and simulation results also underscore the importance of cell-cell signaling in maintaining reproducible lineage proportions at a global scale while facilitating correct pattern formation locally [46, 113, 106].
@@
We find that system size (number of cells) does not significantly influence the accuracy of attaining the correct lineage proportions. This is inline with previous studies reporting that the mouse blastocyst exhibits resilience to ICM size variations, maintaining consistent patterning irrespective of cell number [93, 113]. Although, conversely, we find the precision of cell specification to be system-size dependent, our results suggest that cell-fate misspecification can be reduced down to when the system size surpasses cells, which is comparable to the typical number of cells in the ICM around E3.5. Moreover, our simulations show that increased variability in initial conditions at the cellular level does not necessarily constitute a detriment for the tissue-level dynamics. Instead, increases in the initial-conditions variability, if not excessive, can enhance the accuracy while sustaining the precision of cell-fate specification.
@@
Our ITWT model also successfully recapitulates the temporal sensitivity of the ICM to exogenous FGF4. We observe a specific time window during which the ICM can respond to external FGF4, with the ability to adjust lineage proportions depending on the timing and dosage of the addition. This finding aligns with recent experimental observations [106] and underscores the importance of timing in developmental processes.
@@
A paramount challenge in biophysical mechanistic modeling is the estimation of parameter values allowing the constructed model to faithfully recapitulate the characteristics of the considered biological system. This task becomes particularly complex for spatial-stochastic and mechanistic models due to the need for analyzing behavior across numerous independent samples, significantly increasing the computational demands for navigating their vast parameter spaces [147]. In response to this challenge, we employed an innovative approach by integrating Simulation-Based Inference (SBI) with traditional machine learning techniques. This strategy utilizes the Sequential Neural Posterior Estimation (SNPE) algorithm [59, 27, 32], leveraging simulation data to efficiently traverse parameter space.
@@
This method not only relies on the direct analysis of simulation outcomes but also incorporates qualitative data to identify parameter distributions that align with the expected behaviors of the system, encoded in high-level, low-dimensional utility functions (“target scores”). We utilized a state-of-the-art SBI toolbox [136], which facilitated the integration of the SNPE algorithm into our workflow. We successfully trained artificial neural networks to predict model parameter sets capable of reproducing the targeted ICM patterning behavior of several model variants, corresponding to both wild-type and mutant systems. This enabled us to thoroughly examine the extensive parameter space of our ICM model.
@@
With this approach, we demonstrate that despite the scarcity of detailed quantitative experimental measurements, empirical qualitative observations provide viable bases for biophysical mechanistic modeling. To our knowledge, our work constitutes the first application of an AI-powered simulation-based inference (SBI) framework to spatial-stochastic modeling in developmental biology.
@@
While here we focused on a minimal spatial geometry for targeted assessment of the interplay between biochemical stochasticity and spatial coupling, ICM development is influenced by important additional factors, such as cell divisions and mechanical interaction among cells. Future elaborations of our framework will incorporate suitable tissue-scale dynamics, which will integrate the stochastic dynamics of single-cell gene expression and inter-cellular signaling with the constant remodeling of the tissue geometry. This will enable the study of how cell neighborhoods varying both in time and space influence ICM lineage differentiation, while exploiting recently recorded tissue structural data [46, 48, 150, 47].
@@
In conclusion, our model underscores the complexity and robustness of ICM lineage specification, generating unique insights into the interplay of stochasticity, intercellular signaling mechanisms, and system size (number of cells) in tissue development. These findings not only deepen our understanding of the developing early mouse embryo but also provide a comprehensive framework for exploring similar controlled stochastic processes in other biological systems.

### 13 Mar 2024 v2

@@ Abstract
Understanding how multicellular organisms reliably orchestrate cell-fate decisions is a central challenge in developmental biology. This is particularly intriguing in early mammalian development, where early cell-lineage differentiation arises from processes that initially appear cell-autonomous but later materialize reliably at the tissue level. In this study, we develop a multi-scale, spatial-stochastic simulator of mouse embryogenesis, focusing on inner-cell mass (ICM) differentiation in the blastocyst stage. Our model features biophysically realistic regulatory interactions and accounts for the innate stochasticity of the biological processes driving cell-fate decisions at the cellular scale. We advance event-driven simulation techniques to incorporate relevant tissue-scale phenomena and integrate them with Simulation-Based Inference (SBI), building on a recent AI-based parameter learning method: the Sequential Neural Posterior Estimation (SNPE) algorithm. Using this framework, we carry out a large-scale Bayesian inferential analysis and determine parameter sets that reproduce the experimentally observed system behavior. We elucidate how autocrine and paracrine feedbacks via the signaling protein FGF4 orchestrate the inherently stochastic expression of fate-specifying genes at the cellular level into reproducible ICM patterning at the tissue scale. This mechanism is remarkably independent of the system size. FGF4 not only ensures correct cell lineage ratios in the ICM, but also enhances its resilience to perturbations. Intriguingly, we find that high variability in intracellular initial conditions does not compromise, but rather can enhance the accuracy and precision of tissue-level dynamics. Our work provides a genuinely spatial-stochastic description of the biochemical processes driving ICM differentiation and the necessary conditions under which it can proceed robustly.
@@ Introduction
In mammalian development, the cell signaling and fate specification dynamics during the preimplantation stage of mouse embryogenesis serve as key catalysts for understanding this core paradigm. These processes have been extensively investigated through both experimental and theoretical approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], and rely on dynamic cross-interactions between gene expression, molecular signaling at the tissue level, and mechanical cues essential for tissue remodeling [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@
The preimplantation stage in mouse embryos involves two pivotal events transforming the zygote into a blastocyst composed of three distinct cell types [3, 33, 25, 24, 34]. Initially, cell segregation results in the formation of the trophectoderm (TE), which contributes to the embryonic part of the placenta, and the inner cell mass (ICM), an uncommitted bipotent progenitor tissue. Subsequently, a second cell-fate decision within the ICM leads to the differentiation of epiblast (EPI) and primitive endoderm (PRE) tissues. The EPI, marked primarily by NANOG expression, is pluripotent and gives rise to all embryo-proper tissues. In contrast, the PRE, marked primarily by GATA6 expression, contributes to the formation of extraembryonic supporting tissues. Notably, EPI cells secrete FGF4, a signaling ligand that promotes PRE cell specification. The FGF/ERK signaling pathway therefore plays a crucial role in regulating the balance between EPI and PRE cells. Interestingly, FGF4 acts within an external feedback loop that ultimately inhibits its own production, downregulating it in PRE cells [35, 12, 23]. The mechanisms underlying the precise and timely distribution of FGF4 across the ICM, critical for appropriate cell lineage specification, remain elusive [36, 37].
@@
For successful embryogenesis, the TE, EPI, and PRE lineages in the blastocyst must be segregated in specific proportions and within a narrow developmental time window [38, 12, 25]. This specification occurs over approximately 48 hours between embryonic day (E) 2.5 and E4.5 [23, 12]. By E3.0, when blastocyst formation begins, ICM cells co-express NANOG and GATA6. In the further course, EPI and PRE cells emerge asynchronously from ICM cells in a spatially stochastic manner, influenced by the short-range FGF4 signal and regulated by auto- and paracrine feedback loops [39, 25, 36]. A mutually exclusive gene expression profile emerges, with EPI cells exhibiting high NANOG and low GATA6 levels, and vice versa in PRE cells.
This process is driven by a gene-regulatory network characterized by self-activation and mutual repression of NANOG and GATA6, following a common motif for generating multistable expression states [40, 41, 42, 43, 44], and is refined by the external FGF4 feedback. The spatio-temporal segregation of EPI and PRE lineages is subsequently coordinated through a mechanical cell-sorting process [32]. The final spatially ordered pattern thus is an emergent property of the differentiating tissue, in stark contrast to systems which employ morphogen gradients for high-precision system-wide coordination of developmental trajectories [45, 46, 47, 48, 49, 50, 51, 52], although layers of interacting genes downstream of the gradients can acquire self-organizing and scaling capabilities after their position-dependent activation [53, 54, 55, 56, 57, 58, 44, 59].
@@
Experimental studies have sought to characterize the dynamics and biophysical parameters of ICM specification, focusing on internal and external developmental cues, molecular mechanisms underpinning gene transcription and mRNA translation, and key signaling pathway components [60, 61, 7, 62, 39]. Nevertheless, this remains challenging, as the involved biochemical species are present at low mRNA and protein abundance, leading to significant biological noise. Here biological noise is categorized as either intrinsic (related to the discrete and stochastic nature of biochemical reactions and molecular transport within cells) or extrinsic (referring to external fluctuations affecting all cells). This noise strongly impacts the reliability of developmental dynamics, particularly during early development [63, 64, 65, 66, 67, 68, 69], and hampers experimental measurements at the same time.
@@
More recently, deep-learning-based approaches have been successfully used for both pattern recognition and reconstruction of ICM organoid data, employing experimental and synthetic datasets for training [70]. Although these approaches have good predictive power for both cell-fate spatial composition and determination, they do not reveal full mechanistic insights into the gene regulatory interactions that orchestrate ICM patterning.
In this study, we investigate cell specification during blastocyst formation in the early mouse embryo under truly stochastic conditions. In order to elucidate mechanisms that enable robust and replicable cell-type proportioning in the ICM, we have constructed a spatial model that allows for exact simulation of the stochastic dynamics of key biochemical species in the developing ICM. This model incorporates three distinct diffusive-signaling modes between neighboring cells: autocrine, paracrine, and intermembrane ligand-exchange (akin to juxtacrine) signaling via FGF4.
@@ Results
In computational modeling of complex biological systems, inferring parameter sets that reproduce experimental observations is a key challenge. This usually requires specific domain knowledge, as the selection of an appropriate parameter search method is chiefly influenced by the particular properties of the problem at hand [73]. One significant hurdle in developmental biology is the lack of a versatile, general-purpose inference method suitable for high-dimensional stochastic models, which typically require large and comprehensive datasets with fine-grained resolution. This expands both in terms of the features measured (e.g., gene expression levels at the single-cell scale rather than bulk measurements) and the temporal resolution (e.g., time-series data capturing dynamics at all relevant scales). However, most experimental studies are unable to simultaneously measure all system variables required for such in-depth mechanistic representation. To bridge this gap, Biologically-Informed Neural Networks (BINNs) and Simulation-Based Inference (SBI) frameworks have emerged as powerful strategies addressing several of these modeling and inferential challenges [74, 75, 76, 77]. Particularly advantageously, SBI allows to learn multidimensional parameter sets that comply with a prescribed behavior formulated in terms of lower-dimensional utility functions, thus mitigating the limitations posed by the scarcity of detailed quantitative data.
@@
Estimate the maximum-a-posteriori (MAP) parameter value that aligns with the desired system behavior represented by the target score time series, using the posterior distribution predicted by the ANN conditioned on the given target score.
@@
For a more detailed description of these steps, see Fig 1 and Model parameter inference framework. In a companion study [83] (in preparation), we compare this SNPE-guided methodology against a strategy inspired by a classical optimization algorithm, specifically simulated annealing; we find that the AI-based approach surpasses its classical counterpart in performance, assuming equal allocation of computational resources. This superiority can be traced back to the innate flexibility of the ANN and the absence of a native parameter interpolation procedure in simulated annealing.
@@
Secondly, blastocyst formation occurs within approximately 1.5-2 days of development, starting around E2.75 [19, 21, 25]. The EPI and PRE populations must reach and sustain the required fate proportions 8 to 12 hours before the end of the preimplantation period (around E4.75). This requirement serves as another optimization criterion for our model, regardless of the actual underlying cell-differentiation mechanism.
@@
Fig 2 summarizes the key differences between the ITWT and RTM systems, contrasting the respective inferred parameter values, i.e. the maximum-a-posteriori-probability (MAP) estimates, for both the central intracellular GRN components (top row) and the extracellular signaling topology (bottom row). The most notable difference between the ITWT and RTM models is the reversal of the relationship between the self-activation thresholds of Nanog and Gata6. For the ITWT, the half-saturation threshold of Nanog self-activation (Nanog_NANOG) is lower/stronger than the half-saturation threshold of Gata6 self-activation (Gata6_GATA6). However, both models exhibit a well-balanced mutual repression between Nanog and Gata6, albeit with varying interaction strengths. We present a complete overview of the full model parameter distributions inferred in this study, and a first approximation of the model parameter sensitivities, in the Supporting information, Figs 14 and 15.
@@
The TM-A model displays a decreased accuracy of the ICM specification mechanism due to the absence of self-regulation, though the overall precision of the system remains unaffected (Fig 6[B, D]). Adjustments in other signaling components could potentially correct for this, but these also would likely reduce the system’s ability to buffer against dynamic signaling perturbations.
@@
A central challenge in developmental biology is the precise characterization of spatial patterns, such as the “salt-and-pepper” arrangement reported for the ICM, which has often been described informally in the literature [33, 88, 6, 1, 96]. The term typically implies a random distribution, yet randomness in a mathematical context can take various forms. Recent studies have endeavored to rigorously define this pattern using experimental and theoretical approaches [16, 30, 97, 70, 98]. Here we understand the “salt-and-pepper” pattern as an archetype in which each individual cell-fate decision is independent of the cell fates of its neighbors, which implies a multinomial distribution of cell fates in every tissue neighborhood.
@@
Our approach was guided by two aspects: firstly, the broad inferred posterior distributions of the parameters governing the ICDs, and their moderate sensitivity to value changes (Fig 15[D, H]); secondly, similar assessments that were carried out in previous models of mouse-blastocyst, where mainly the ICD variance was modulated [5, 13]. Taking this into account, we generated 1000 stochastic trajectories for 10 distinct ICDs, modifying their variance from 0% to 200% compared with the baseline value (see Fig 8[A] for examples); the details of ICD modulation are described in Computational experiments.
We analyzed both the mean cell count (accuracy) and the corresponding standard deviation (precision) for each cell fate (Fig 8[C, D]).
@@
We quantified the deviations from the typical ICD behavior using the relative Fate Error (FE), which measures the discrepancy in ICM specification accuracy for each cell fate (UND/EPI/PRE) between the perturbed scenarios and the reference (0% ICP) scenario at 48 hours; see also caption of Fig 9 for details about FE.
@@
Temporal modulation of FGF4 concentration and the corresponding shift in cell plasticity are key aspects of ICM cell differentiation, with numerous studies documenting their influence [12, 107, 36, 92, 25]. To examine whether our model can replicate the experimentally observed time-dependent responsiveness to FGF4 level changes, we introduced controlled perturbations of FGF4 in our simulations. This involved adding extra FGF4 molecules to each simulated cell at specific time points, mimicking the effect of exogenous FGF4.
We assessed the response to FGF4 perturbations in two models (ITWT and TM-FGF4) and at two distinct final simulation times, 48 and 96 hours. The second time point, while biologically irrelevant, was introduced to test whether manipulation of FGF4 levels can alter the typical time scale on which cell-fate commitment converges. Exogenous FGF4 addition was carried out at predetermined time points: hours, with the amount of added FGF4 molecules determined by a mix of Poissonian and binomial distributions (details in Computational experiments).
Note that the TM-FGF4 system represents a theoretical mutant with blocked FGF4 production, mimicking a full loss-of-function phenotype for the Fgf4 gene; this means that any FGF4 originates in these systems from the external addition.
We quantified the ICM patterning robustness to varying amounts of exogenous FGF4 using the Standardized Absolute Fate Deviation (SAFD), which measures the distance, in terms of mean cell-fate proportions, from the null configuration (no exogenous FGF4 at 0 hours) to every alternative perturbation-magnitude-and-addition-time configuration; see also caption of Fig 10 for details about SAFD.
Our analysis at the 48-hour mark shows that the TM-FGF4 model is highly responsive to exogenous FGF4 added at the start of the simulation (addition time = 0 hours). Depending on the FGF4 count, it is possible to manipulate the lineage ratios and rescue the PRE fate (Fig 10[E]), which is in line with recent in vitro experiments [36].
However, we identify an end-of-plasticity time point around 24 hours, after which the system becomes insensitive to additional FGF4 and locks into its pre-existing cell lineage proportions (Fig 10[F-H]). In the ITWT system, cell fate ratios can also be manipulated by varying the FGF4 amount, although no distinct end-of-plasticity time point is observed in this model (Fig 10[A-D]).
@@ Discussion
Several phenomenological models have been proposed for blastocyst formation in the mouse embryo [33, 5, 23, 13], but they are primarily deterministic in nature, and as such do not allow for a rigorous investigation of the implications of noise emerging from the basic processes driving this developmental process, neither for quantification of its robustness. To correctly capture the inherent randomness in ICM differentiation, we developed a biophysics-rooted spatial-stochastic model simulated via the Reaction-Diffusion Master Equation (RDME) formalism, and embedded it into a Simulation-Based Inference (SBI) framework building on recent advancements in Machine Learning (ML). Our multi-cellular model mechanistically describes the biochemical ICM patterning dynamics and its accuracy using biologically realistic lifetimes for the involved molecular species, and provides a biophysically realistic implementation of the mesoscopic processes generating noise at the cell level. Using this combined framework, we inferred multiple parameter distributions that inform our model both in wild-type-like and several mutant-like conditions.
@@
Our inferred theoretical wild-type (ITWT) model recovers key experimental findings, specifically: high reproducibility of the EPI and PRE lineage proportions, the observed timescale of blastocyst formation, amounting to 1.5-2 days of embryonic development, and the indispensability of FGF4-mediated signaling for appropriate ICM patterning [33, 84, 25]. In the absence of FGF4 coupling, the simulated blastocyst fails to establish correct cell fate proportions, displaying a strong bias towards the EPI fate.
We thus argue that, given a default naive pluripotent state in ICM-like systems, successful cell-fate specification necessitates a tissue-level coordinating mechanism. Previous experimental and simulation results also underscore the importance of cell-cell signaling in maintaining reproducible lineage proportions globally while facilitating correct pattern formation locally [30, 23, 36]. Interestingly, our analysis of cell neighborhood composition reveals that the communication range of FGF4, though essentially limited to nearest neighbors, suffices for ensuring effective signaling. This observation aligns with recent findings from both in vitro and in silico studies [36, 70], highlighting the nuanced roles of local signaling dynamics in complex tissue patterning.
@@
We find that system size (number of cells) does not significantly influence the accuracy of attaining the correct lineage proportions. This is in line with previous studies reporting that the mouse blastocyst exhibits resilience to ICM size variations, maintaining consistent patterning irrespective of cell number [11, 23]. Although, conversely, we find the precision of cell specification to be system-size dependent, our results suggest that cell-fate misspecification can be reduced down to when the system size surpasses cells, which is comparable to the typical number of cells in the ICM around E3.5. Moreover, our simulations show that increased variability in initial conditions at the cellular level does not necessarily constitute a detriment for the tissue-level dynamics. Instead, increased variability in initial molecular resources, if not excessive, can enhance the accuracy while sustaining the precision of cell-fate specification.
@@
Our ITWT model also successfully recapitulates the temporal sensitivity of the ICM to exogenous FGF4. We observe a specific time window during which the ICM can respond to external FGF4, with the ability to adjust lineage proportions depending on the timing and dosage of the addition. This finding aligns with recent experimental observations [36] and highlights the importance of timing in developmental processes.
@@
A paramount challenge in biophysical mechanistic modeling is the estimation of parameter values allowing the constructed model to faithfully recapitulate the characteristics of the considered biological system. This task becomes particularly complex for spatial-stochastic and mechanistic models due to the need for analyzing behavior across numerous independent simulation samples, significantly increasing the computational demands for navigating their vast parameter spaces [108, 77].
In response to this challenge, our approach integrates an AI-powered Simulation-Based Inference (SBI) method with traditional machine learning techniques, specifically employing the Sequential Neural Posterior Estimation (SNPE) algorithm [71, 74, 72]. This strategy leverages simulation data to efficiently traverse parameter space, incorporating both direct analysis of simulation outcomes and qualitative data to identify parameter distributions that align with the expected behaviors of the system, encoded in high-level, low-dimensional utility functions (“target scores”).
@@
We utilized a state-of-the-art SBI toolbox [109], which facilitated the integration of the SNPE algorithm into our workflow. This allowed us to train artificial neural networks (ANNs) for predicting model parameter sets capable of reproducing the targeted ICM patterning behavior of several model variants, corresponding to both wild-type and mutant systems.
Our methodology imposes minimal constraints on the inference problem by leveraging only essential experimental observations. This strategy prevents model overfitting, allowing for the extrapolation of system behaviours spanning multiple spatial and temporal scales not directly observed in experimental data.
@@
With this approach, we underscore the distinctions between machine learning models, which prioritize universal prediction at the expense of modeling interpretability, and mechanistic models, which focus on exploratory hypotheses to uncover causal relationships at the expense of modeling fidelity [73, 110]. By merging these paradigms, we demonstrate that despite the scarcity of detailed quantitative experimental measurements, the flexibility and predictive capabilities of ANNs can aid generating full-featured quantitative predictions by imposing key empirical qualitative observations to mechanistic biophysical models. To our knowledge, our work constitutes the first application of an AI-powered SBI framework to spatial-stochastic modeling in developmental biology.
@@
While here we focused on a minimal spatial geometry for targeted assessment of the interplay between biochemical stochasticity and spatial coupling, ICM development is influenced by important additional factors, such as cell divisions and mechanical interaction among cells. Future elaborations of our framework will incorporate suitable tissue-scale dynamics, which will integrate the stochastic dynamics of single-cell gene expression and inter-cellular signaling with the constant remodeling of the tissue geometry.
This will enable the study of how cell neighborhoods varying both in time and space influence ICM lineage differentiation, while exploiting recently recorded tissue structural data [30, 97, 32, 98].
@@
In conclusion, our AI-parameterized model underscores the complexity and robustness of the EPI-PRE lineage specification, generating unique insights into the interplay of stochasticity, biochemical feedbacks, tissue-level signaling mechanisms, and system size (number of cells) in ICM development. These findings not only deepen our understanding of the developing early mouse embryo but also provide a comprehensive framework for exploring similar controlled stochastic processes in related biological systems.